# MCGAN: Enhancing GAN Training with Regression-Based Generator Loss

### v1

@@ 4.3 Relation to feature matching
5 Numerical experiments
To validate the effectiveness of the proposed MCGAN method, we conduct extensive experiments on a variety of generation tasks on image, time-series, and video data. We mainly focuses on the images for in-depth analysis in this section, and briefly introduce our attempts for the extensions to time-series and video data in the last two subsections.
The training algorithm of MCGAN for conditional generation task can be found in Algorithm 1 in Appendix C. We also put extra experimental details in Appendix B for reproducibility.
5.1 Datasets for image generation
[…]
5.2 Architectures, hyperparameters, and training techniques
5.2.1 Backbone GAN architectures
We employ BigGAN [27] and cStyleGAN2 [28] architectures as the backbones for image generation experiments.
BigGAN [27], as a member of projection-based cGAN, is a collection of recent best practices in conditional image generation, and it is widely used due to its satisfactory generation performance on high-fidelity image synthesis. We use the BigGAN architecture with the same regularization methods like Exponential Moving Averages (EMA) [29] and Spectral Normalization [4] have already been adopted. We adopt BigGAN’s PyTorch implementation  [^3] https://github.com/PeterouZh/Omni-GAN-PyTorch and shows the architectural details in Table 5 for completeness.
cStyleGAN2 [30], is an improved and conditional version of the original StyleGAN, is a generative adversarial network (GAN) architecture designed for creating high-quality, diverse images. It addresses artifacts, enhances image quality, and has been widely used for generating realistic portraits and artwork. We adopt the code from the Github repository in [30]  [^4] https://github.com/mit-han-lab/data-efficient-gans/tree/master. and use the default hyperparameter setting. The only difference is the minor modification we made to incorporate our MC method.
5.2.2 Loss functions
Since our MCGAN replaces the original generative loss in Eqn. (4) with the regression loss in Eqn. (5) and keeps the discriminator loss, we use two popular discriminator’s loss functions as baselines: the Hinge loss baseline and the BCE loss baseline.
5.2.3 Hyperparameters
For the CIFAR-10 experiment, the batch size is set to 32. We adopt the Adam optimizer in all experiments, with betas being 0.0 and 0.999. For both of the generator and discriminator, the learning rates are set to 0.0002 and the weight decay is 0.0001. The model is updated by using Exponential Moving Average starting after the first 5000 iterations. The generator is updated once every 3 times the discriminator is updated. For the CIFAR-100 experiment, we have fewer training samples for each class, so we update the discriminator 4 times per generator training as a more accurate discriminator is needed. Each experiment is conducted on one Quadro RTX 8000 GPU.
5.2.4 Training techniques
Leaky Clamp
where $\alpha\in(0,1)$ is a small slope for values outside the range $[lb,ub]$. Just similar to the negative slope in Leaky ReLU [31], the $\alpha$ in the Leaky Clamp function is used to prevent from vanishing gradient problem. And by applying this Leaky Clamp on the discriminator output when computing the regression loss, we are able to mitigate the early collapse problem.
Data augmentation
To alleviate the overfitting and improve the generalization on the small training set, especially for CIFAR-100 where each class has scarce samples, we increase data efficiency by using the Differentiable Augmentation (DiffAug) [30] which imposes various types of augmentations on real and fake samples [30, 32, 33]. We adopt Translation + Cutout policy as suggested in [30]. Besides, we also apply horizontal flips when loading the training dataset as in [13].
5.3 Evaluation metrics
Inception Score [25] (IS) is a popular metric to evaluate the variety and distinctness of the generated images. It is given by
$\text{IS}=\exp\{\mathbb{E}_{X\sim\nu_{\theta}}\left[D_{\text{KL}}(P(Y|X)||P(Y))\right]\}$ (18)
where $D_{\text{KL}}$ is the KL-divergence between the conditional class distribution $P(Y|X)$ and marginal class distribution of the $P(Y)=\mathbb{E}[P(Y|\mathbf{X})]$. The conditional class distribution $P(Y|X)$ is computed by InceptionV3 network pre-trained on ImageNet. The higher IS, the better the quality. By the definition, the IS does not consider real images, so cannot measure how well the fake measure induced by the generator is close to the real distribution. Other limitations, as noted in [34], are: high sensitivity to small changes in weights of the Inception network, and large variance of scores. To consider both diversity and realism, the following FID and IFID are employed as well.
Fréchet Inception Distance [35] (FID) compares the distributions of Inception embeddings of real and generated images, denoted by $p_{d}$ and $p_{\theta}$ respectively. Under the assumption that the features of images extracted by the function $f$ are of multivariate normal distribution. The FID score of $p_{\theta}$ w.r.t $p_{d}$ is defined as
$\text{FID}(p_{d},p_{\theta})=\|\mathbf{\mu}_{r}-\mathbf{\mu}_{g}\|^{2}+\text{Tr}(\mathbf{\Sigma}_{r}+\mathbf{\Sigma}_{g}-2(\mathbf{\Sigma}_{r}\mathbf{\Sigma}_{g})^{\frac{1}{2}}),$ (19)
$\mathbf{\mu}=\frac{1}{N}\sum^{N}_{i=0}f(x_{i}),\quad\Sigma=\frac{1}{N-1}\sum_{i=0}^{N}(f(x^{(i)}-\mathbb{\mu}))(f(x^{(i)}-\mathbb{\mu}))^{T}.$
Intra Fréchet Inception Distance [36] (IFID) is used to quantify intra-class diversity. It is defined as the average of conditional FID given every class $y\in\mathcal{Y}$, i.e.,
where
$\text{FID}(p_{d}(y),p_{\theta}(y))=$
$\|\mathbf{\mu}_{r}(y)-\mathbf{\mu}_{g}(y)\|^{2}+\text{Tr}(\mathbf{\Sigma}_{r}(y)+\mathbf{\Sigma}_{g}(y)-2(\mathbf{\Sigma}_{r}(y)\mathbf{\Sigma}_{g}(y))^{\frac{1}{2}}).$
Recognizability is as crucial as realism and diversity in a good image generative model, therefore two classification accuracy are adopted—a weak accuracy (WA) measured by a two-layer convolutional neural network  [^5] https://pytorch.org/tutorials/beginner/blitz/cifar10_tutorial.html and a strong accuracy (SA) by the ResNet-50 [37]. Both classifiers are pre-trained on the same training set as for the generative model. The WA discerns subtle differences for better inter-model comparison, while the SA is more accurate for intra-model latent space analysis.
5.4 Evaluations of conditional image generation on CIFAR-10
5.4.1 Learning curves in training
5.4.2 Quantitative results
The quantitative results in Table 2 show that our MC method considerably improves all the baselines independently of different discriminative losses. Specifically, when using Hinge loss as the discriminative loss function along with DiffAug, the MC method can improve the FID from 4.43 to 3.61, which is comparable to most state-of-the-art models [5]. Also, its IS score also increased from 9.61 to 9.96, indicating better diversity of the generated samples.
Loss Hinge BCE
Original 9.27 $\pm$ 0.11 5.31 16.20 9.301 $\pm$ 0.14 5.55 16.62
+DiffAug 9.61 $\pm$ 0.06 4.43 14.60 9.51 $\pm$ 0.11 4.71 14.83
+MC 9.66 $\pm$ 0.09 4.51 14.71 9.62 $\pm$ 0.09 4.61 14.82
+MC+DiffAug 9.96 $\pm$ 0.12 3.61 13.60 9.94 $\pm$ 0.10 3.93 13.72
Table 2: Quantitative results of image generation on CIFAR-10 using BigGAN w/o and with our MC method and Differentiable Augmentation.
(a) FID ↓
(b) IS ↑
For the recognizability metrics, we generated 10k (same setting as the test set) images using BigGAN backbone. The WA rates are 62.56%, 52.09%, and 54.71% for the real test set, the generated set from Hinge baseline, and the generated set from Hinge + MC, respectively. Our MC method’s images perform closer to the real test set than the baseline’s images, showing better distribution matching to the real data in terms of recognizability. The SA rate of our MC method is 83.42% compared to 93.65% of the real test set, showing that we generate fairly recognizable fake images.
5.4.3 Qualitative results
The qualitative results are shown in Figure 4 with only a small amount of images (in red boxes) misclassified by our classifier.
5.4.4 Results based on cStyleGAN2 backbone
In addition, we use cStyleGAN2 as the backbone for evaluation, summarize the results in Table 3, and show the qualitative results in Figure 5. We can see from Table 3 that under both Hinge and BCE losses, the proposed MC method is able to improve their FID by around 0.08. Among them, Hinge + MC + DiffAug achieves the best FID 2.16, outperforming most of the algorithms using StyleGAN2 as the backbone [13, 5, 38].
Based on our best model using cStyleGAN2, we plot the FID of each class in Figure 6. It shows that among all classes, the MC method consistently outperforms its counterparts. Cats and dogs are the two classes where all models have their worst FID. The proposed MC method is able to improve the FID of these two classes by around 0.45, which is a considerable gap.
Loss Hinge BCE
+DiffAug 10.1859 2.2475 11.4037 10.1281 2.4387 11.6191
+MC+DiffAug 10.2608 2.1650 11.0351 10.0995 2.3595 11.3004
Table 3: Quantitative results of image generation on CIFAR-10 using cStyleGAN2 w/o and with our MC method and Differentiable Augmentation.
Loss Hinge BCE
Original 10.73$\pm$ 0.10 8.31 83.36 10.81$\pm$ 0.14 8.37 81.89
+DiffAug 10.72$\pm$ 0.13 7.37 80.00 10.71$\pm$ 0.08 7.61 80.48
+MC 11.39$\pm$ 0.10 6.97 80.20 11.59$\pm$ 0.12 6.99 80.91
+MC+DiffAug 11.81$\pm$ 0.06 5.77 76.26 11.90$\pm$ 0.08 5.85 77.33
Figure 6: FID ↓ of each class on CIFAR-10 dataset.
(a) FID
(b) $\mathbb{E}_{\mu}\|\nabla_{x}D^{\phi}(x)\|^{2}_{2}$
(c) $\text{Var}_{\mu}\|\nabla_{x}D^{\phi}(x)\|^{2}_{2}$
Figure 7: The trend of (a) Fréchet Inception Distance; (b) $\mathbb{E}\|\nabla_{x}D^{\phi}(x)\|^{2}_{2}$; (c) $\text{Var}\|\nabla_{x}D^{\phi}(x)\|^{2}_{2}$ for models trained by different methods using CIAFR10 dataset. Here Hinge stands for the model trained by Hinge loss, and Hinge+MC for the model trained by Hinge loss and our proposed MC method. Lower $\text{Var}_{\mu}\|\nabla_{x}D^{\phi}(x)\|^{2}_{2}$ indicate better stability.
5.4.5 Evaluation of training stability
In Figure 7, we also plot the training dynamics of models trained on CIFAR-10 dataset. It is clear that the FID dynamic of the model trained with MC method is more stable than that of the original Hinge GAN and converged to a better level. In Figure 7(b) and 7(c), the dynamics of $\mathbb{E}\|\nabla_{x}D^{\phi}(x)\|^{2}_{2}$ and $\text{Var}\|\nabla_{x}D^{\phi}(x)\|^{2}_{2}$ also indicates that the discriminator gradient $\nabla_{x}D^{\phi}$ of Hinge+MC is less volatile, which can result in better training stability of its generator due to its reliance on the sensitivity of $\nabla_{x}D^{\phi}$
5.4.6 Latent space analysis
The latent space learned by the generator is expected to be continuous and smooth so that small perturbations on the conditional input can lead to smooth and meaningful modifications on the generated output. To explore the latent space, we interpolate between each pair of randomly generated images by linearly interpolating their conditional inputs. The results are shown in Figure 8. Intermediary images between a pair of images from two different classes are shown in each row with their confidence score distributions below. The labels of the two classes are shown on the left and right sides of each row. Each distribution of the confidence scores are calculated by the bottleneck representation of the ResNet-50 classifier with a softened softmax function of temperature 5.0 for normalization. The score bars of the left class and the right class are shown in green and magenta, respectively. The red boxes highlight the images being classified as a third class, while the yellow boxes highlight the images having non-monotonic transitions of their confidence scores compared to those of their adjacent images. In other words, images in both red and yellow boxes are undesirable as they imply that the latent space is less continuous and less smooth. By comparing Figure 8a and 8b, we can see that the MC method performs better in the learned latent space and has most of the decision switch between the two classes occur in the middle range of the interpolation.
5.5 Evaluations of image generation on CIFAR-100
For completeness, we show the image generation performance on CIFAR-100 in Table 4. Significant improvements are achieved by using our MC method independently for both baseline discriminative losses, with average improvement of 1.1 in IS, 1.6 in FID and 3.7 in IFID.
5.6 Extension to conditional time-series generation
Our MCGAN framework is so flexible that can be incorporated into the conditional time-series generation task, in which case, our goal is to generate future paths based on the past path of the time series. To benchmark with MCGAN, we choose three representative generative models for the time-series generation, i.e. (1) TimeGAN [15]; (2) RCGAN [39] - a conditional GAN; and (3) GMMN [40] - an unconditional maximum mean discrepancy (MMD) with a Gaussian kernel. We trained these models on both vector auto-regressive (VAR) data and stock (S&\&&P 500 and DJI) data. The model performance is evaluated using metrics in [41] including (1) ABS metric - a histogram-based metrics that measure the distance between two distributions; (2) ACF metric - a metric that measures the difference in temporal dependency; (3) $R^{2}$ relative error - a metric that demonstrates the effectiveness of synthetic data to be used for real applications. The numerical results illustrate that our MC method can considerably outperform these benchmarks. To be specific, in SPX data, the ABS metric and ACF metric of RCGAN can be improved from 0.01134 to 0.00915 and from 0.03665 to 0.04192 by using our proposed MC method. In VAR dataset, MCGAN also dominates in terms of almost all test metrics under different parameter settings of VAR data.
5.7 Extension to conditional video generation
Inspired by the effectiveness of MCGAN in generation tasks on both spatial image data and temporal time-series data, we also validate our MCGAN model in conditional video generation task. In this task, the generator aims to generate future video frame given the past frames. Here we used Moving MNIST data set [42], which consists of 10,000 20-frame 64x64 videos of moving digits. The whole dataset is divided into the training set (9,000 samples) and the test set (1,000 samples). For the architecture of both generator and discriminator, we use convolutional LSTM (ConvLSTM) unit proposed by [43] due to its effectiveness in video prediction task. In the model training, the generator takes in 5 past frames as the input and generates the corresponding 1-step future frame, then the real past frames and the generated future frames are concatenated along time dimension and put into the discriminator. For comparison, we used classical GAN, which uses BCE as the generative loss, as the benchmark. We trained our model for 10,000 epochs with batch size 16. The model performance is evaluated by computing the MSE between the generated frames and the corresponding ground truth on the test set. From the numerical results, we find that, by using our proposed MC method, the MSE can be improved from 0.153 (GAN) to 0.123 (MCGAN). Compared to the baseline, the predicted frames from our MC method are clearer and more coherent, and visually closer to the ground truth as shown in Figure 9.
Figure 9: Results of predicting future frame given the past 5 frames based on ConvLSTM w/o and with our MC method.
6 Conclusion
This paper presents a general MCGAN method to tackle the training instability, a key bottleneck of GANs. Our method enhances generator training by introducing a novel regression loss for (conditional) generative adversarial networks. We establish the optimality and discriminativity of MCGAN, and prove that the convergence of optimal generator can be achieved under a weaker condition of the discriminator due to the strong supervision of the regression loss. Moreover, extensive numerical results on image data such as Cifar 10 and 100, time series data, and video data are provided to validate the effectiveness and flexibility of our proposed MC-GAN and consistent improvements over the benchmarking GAN models.
For future work, we plan to explore the potential of the MCGAN on larger datasets for high-resolution image generation. We also plan to investigate the effectiveness of the MCGAN in more complex and challenging generation tasks, such as text-to-image generation and image-to-image translation. Last but not least, given the flexibility of the MCGAN on different types of data, we consider generation tasks on data with multiple modalities at the same time.
Appendix A Proofs
A.1 Proof of optimality
Proof  For every $\phi\in\Phi$, the derivative of $L_{G}(\theta;\phi,\mu)$ in (5) w.r.t $\theta$ can be derived as
@@ A.1 Proof of optimality
By Assumption 3.1, equation (21) holds only if $\nu_{\theta^{*}}=\mu$. Here we focus on the other case (20).
@@
for every different $\mu$ and $\nu_{\theta}$. Hence equality (22) holds if and only if $\nu_{\theta^{*}}=\mu,$ which completes the proof.
A.2 Proof of $f$-divergence
See 4.1
Proof  Given the optimal discriminator in (13), we have
@@ A.2 Proof of $f$-divergence
which completes the proof.
A.3 Proof of improved stability
See 4.1
Proof  Since $D^{\phi_{\epsilon}}(x)=D^{\phi^{*}}(x)+\epsilon_{1}(x)$ and $\nabla_{x}D^{\phi_{\epsilon}}(x)=\nabla_{x}D^{\phi^{*}}(x)+\epsilon_{2}(x)$, we have
@@ A.3 Proof of improved stability
$=$ $\left(\mathbb{E}_{\nu_{\theta}}[D^{\phi^{*}}(x)]-\mathbb{E}_{\mu}[D^{\phi^{*}}(x)]+\mathbb{E}_{\nu_{\theta}}[\epsilon_{1}(x)]-\mathbb{E}_{\mu}[\epsilon_{1}(x)]\right)H(\theta,\phi_{\epsilon})$
@@
$\nabla_{\theta}L_{G}(\theta;\phi_{\epsilon})=$ $\Delta(\theta,\phi^{*})H(\theta,\phi^{*})+\bar{\epsilon}_{1}(\theta)H(\theta,\phi^{*})+\Delta(\theta,\phi^{*})\bar{\epsilon}_{2}(\theta)+\bar{\epsilon}_{1}(\theta)\bar{\epsilon}_{2}(\theta)$
$=$ $\nabla_{\theta}L_{G}(\theta;\phi^{*})+\bar{\epsilon}_{1}(\theta)H(\theta,\phi^{*})+\Delta(\theta,\phi^{*})\bar{\epsilon}_{2}(\theta)+\bar{\epsilon}_{1}(\theta)\bar{\epsilon}_{2}(\theta).$
Since both $\bar{\epsilon}_{1}(\theta)$ and $\bar{\epsilon}_{2}(\theta)$ are weighted averages or linear combinations of centred random noises, they are both centred noises as well. Moreover, the expectation of $\bar{\epsilon}_{1}(\theta)\bar{\epsilon}_{2}(\theta)$ is also zero since $\epsilon_{1}(x)$ and $\epsilon_{2}(x)$ are uncorrelated. Hence the mean of $\nabla_{\theta}L_{G}(\theta;\phi_{\epsilon})$ equals to $\nabla_{\theta}L_{G}(\theta;\phi^{*}).$ By the definition of $\Delta(\theta,\phi)$ and $\bar{\epsilon}_{1}(\theta)$, its variance also depends on the difference between $\mu$ and $\nu_{\theta}$, which completes the proof.
A.4 Discriminability
@@
  • Vanilla GAN [1]: GAN employs BCE as the discriminative loss function defined as
    $L_{D}(\phi;\mu,\nu_{\theta})=\mathbb{E}_{\mu}\left[\log(D^{\phi}(X))\right]+\mathbb{E}_{\nu_{\theta}}\left[\log(1-D^{\phi}(X))\right].$ (24)
    As proven in [1], the optimal discriminator given binary cross-entropy loss can be derived as:
@@
  • Least Square GAN [8]: LSGAN employs least square loss function defined as follows:
    $L_{D}(\phi;\mu,\nu_{\theta})=-\mathbb{E}_{\mu}\left[(D^{\phi}(X)-\alpha)^{2}\right]-\mathbb{E}_{\nu_{\theta}}\left[(D^{\phi}(X)-\beta)^{2}\right],$
@@
  • Geometric GAN [7]: Hinge loss function is defined as
    $L_{D}(\phi;\mu,\nu_{\theta})=-\mathbb{E}_{\mu}\left[\max(0,1-D^{\phi}(X))\right]-\mathbb{E}_{\nu_{\theta}}\left[\max(0,1+D^{\phi}(X))\right].$
    By Lemma B.1 in [7], it is straightforward to show that the optimal discriminator can be derived as:
    $D^{\phi^{*}_{\mu,\nu_{\theta}}}(x)=2\mathbbm{1}_{\{p_{\mu}(x)\geq p_{\nu_{\theta}}(x)\}}-1.$ (27)
    It is clear that $D^{\phi^{*}}(x)=f(\frac{p_{\mu}(x)}{p_{\nu_{\theta}}(x)})$, where $f(l)=2\mathbbm{1}_{\{l\geq 1\}}-1$, and $D^{\phi^{*}}(x)(p_{\mu}(x)-p_{\nu_{\theta}}(x))>0$ when $p_{\mu}(x)\neq p_{\nu_{\theta}}(x).$
  • Energy-based GAN [23]: Energy-based loss function is defined as
    $L_{D}(\phi;\mu,\nu_{\theta})=-\mathbb{E}_{\mu}\left[D^{\phi}(X)\right]-\mathbb{E}_{\nu_{\theta}}\left[\max(0,m-D^{\phi}(X))\right].$
    where $m>0$. By Lemma 1 in [23], the optimal discriminator given energy-based loss function can be derived as:
    $D^{\phi^{*}_{\mu,\nu_{\theta}}}(x)=m\mathbbm{1}_{\{p_{\mu}(x)<p_{\nu_{\theta}}(x)\}}.$ (28)
@@
  • $f$-GAN [44]: In $f$-GAN, variational lower bound (VLB) on the $f$-divergence $\text{Div}_{f}(\mu||\nu_{\theta})$ is used in the generative-adversarial approach to mimic the target distribution $\nu_{\theta}$. Let $f:\mathbb{R}_{+}\to\mathbb{R}$ be a convex, lower-semicontinuous function. In $f$-GAN, the discriminative loss is defined as the variational lower bound on certain $f$-divergence:
    $L_{D}(\phi;\mu,\nu_{\theta})=\mathbb{E}_{\mu}\left[D^{\phi}(X)\right]-\mathbb{E}_{\nu_{\theta}}\left[f^{*}(D^{\phi}(X))\right],$ (29)
    where $f^{*}$ is the convex conjugate function of $f$. Under mild conditions on function $f$ [45], the maximum of (29) is achieved when

### v3

@@ Appendix B Properties of MCGAN
Appendix C Proofs
Proof of optimality in Theorem 1
Proof.
For every $\phi\in\Phi$, the derivative of $L_{G}(\theta;\phi,\mu)$ in (3) w.r.t $\theta$ can be derived as
@@ Appendix C Proofs
By Assumption 1, equation (15) holds only if $\nu_{\theta^{*}}=\mu$. Here we focus on the other case (14).
@@
for every different $\mu$ and $\nu_{\theta}$. Hence equality (16) holds if and only if $\nu_{\theta^{*}}=\mu,$ which completes the proof. ∎
Proof of $f$-divergence in Lemma 1
Proof.
Given the optimal discriminator in (10), we have
@@
which completes the proof. ∎
Proof of improved stability in Theorem 2
Proof.
Since $D^{\phi_{\epsilon}}(x)=D^{\phi^{*}}(x)+\epsilon_{1}(x)$ and $\nabla_{x}D^{\phi_{\epsilon}}(x)=\nabla_{x}D^{\phi^{*}}(x)+\epsilon_{2}(x)$, we have
@@
$\nabla_{\theta}L_{G}(\theta;\phi_{\epsilon})=$ $\nabla_{\theta}L_{G}(\theta;\phi^{*})+\bar{\epsilon}_{1}(\theta)H(\theta,\phi^{*})$
$+\Delta(\theta,\phi^{*})\bar{\epsilon}_{2}(\theta)+\bar{\epsilon}_{1}(\theta)\bar{\epsilon}_{2}(\theta).$
Since both $\bar{\epsilon}_{1}(\theta)$ and $\bar{\epsilon}_{2}(\theta)$ are weighted averages or linear combinations of centered random noises, they are both centered noises as well. Moreover, the expectation of $\bar{\epsilon}_{1}(\theta)\bar{\epsilon}_{2}(\theta)$ is also zero since $\epsilon_{1}(x)$ and $\epsilon_{2}(x)$ are uncorrelated. Hence the mean of $\nabla_{\theta}L_{G}(\theta;\phi_{\epsilon})$ equals to $\nabla_{\theta}L_{G}(\theta;\phi^{*}).$ By the definition of $\Delta(\theta,\phi)$ and $\bar{\epsilon}_{1}(\theta)$, its variance also depends on the difference between $\mu$ and $\nu_{\theta}$, which completes the proof. ∎
Discriminability of different GAN variants
@@
  • Vanilla GAN (Goodfellow et al. 2014): GAN employs BCE as the discriminative loss function defined as
    $L_{D}(\phi;\mu,\nu_{\theta})=$ $\mathbb{E}_{\mu}\left[\log(D^{\phi}(X))\right]$ (18)
    $+\mathbb{E}_{\nu_{\theta}}\left[\log(1-D^{\phi}(X))\right].$
    As proven in (Goodfellow et al. 2014), the optimal discriminator given binary cross-entropy loss can be derived as:
@@
  • Least Square GAN (Mao et al. 2017): LSGAN employs least square loss function defined as follows:
    $L_{D}(\phi;\mu,\nu_{\theta})=$ $-\mathbb{E}_{\mu}\left[(D^{\phi}(X)-\alpha)^{2}\right]$
    $-\mathbb{E}_{\nu_{\theta}}\left[(D^{\phi}(X)-\beta)^{2}\right],$
@@
  • Geometric GAN (Lim and Ye 2017): Hinge loss function is defined as
    $L_{D}(\phi;\mu,\nu_{\theta})=$ $-\mathbb{E}_{\mu}\left[\max(0,1-D^{\phi}(X))\right]$
    $-\mathbb{E}_{\nu_{\theta}}\left[\max(0,1+D^{\phi}(X))\right].$
    By Lemma B.1 in (Lim and Ye 2017), it is straightforward to show that the optimal discriminator can be derived as:
    $D^{\phi^{*}_{\mu,\nu_{\theta}}}(x)=2\mathds{1}_{\{p_{\mu}(x)\geq p_{\nu_{\theta}}(x)\}}-1.$ (21)
    It is clear that $D^{\phi^{*}}(x)=f(\frac{p_{\mu}(x)}{p_{\nu_{\theta}}(x)})$, where $f(l)=2\mathds{1}_{\{l\geq 1\}}-1$, and $D^{\phi^{*}}(x)(p_{\mu}(x)-p_{\nu_{\theta}}(x))>0$ when $p_{\mu}(x)\neq p_{\nu_{\theta}}(x).$
  • Energy-based GAN (Zhao 2016): Energy-based loss function is defined as
    $L_{D}(\phi;\mu,\nu_{\theta})=$ $-\mathbb{E}_{\mu}\left[D^{\phi}(X)\right]$
    $-\mathbb{E}_{\nu_{\theta}}\left[\max(0,m-D^{\phi}(X))\right].$
    where $m>0$. By Lemma 1 in (Zhao 2016), the optimal discriminator given energy-based loss function can be derived as:
    $D^{\phi^{*}_{\mu,\nu_{\theta}}}(x)=m\mathds{1}_{\{p_{\mu}(x)<p_{\nu_{\theta}}(x)\}}.$ (22)
@@
  • $f$-GAN (Nowozin, Cseke, and Tomioka 2016): In $f$-GAN, variational lower bound (VLB) on the $f$-divergence $\text{Div}_{f}(\mu||\nu_{\theta})$ is used in the generative-adversarial approach to mimic the target distribution $\nu_{\theta}$. Let $f:\mathbb{R}_{+}\to\mathbb{R}$ be a convex, lower-semicontinuous function. In $f$-GAN, the discriminative loss is defined as the variational lower bound on certain $f$-divergence:
    $L_{D}(\phi;\mu,\nu_{\theta})=$ $\mathbb{E}_{\mu}\left[D^{\phi}(X)\right]$ (23)
    $-\mathbb{E}_{\nu_{\theta}}\left[f^{*}(D^{\phi}(X))\right],$
    where $f^{*}$ is the convex conjugate function of $f$. Under mild conditions on function $f$ (Nguyen, Wainwright, and Jordan 2010), the maximum of (23) is achieved when